OpenStudy (anonymous):

joshua and ashley buy three whole pizza. joshua ate 1 1/4 of a pizzas and ashley ate 1 1/5 of a pizza what fraction did they take home

OpenStudy (anonymous):

first find how much pizza they ate.

OpenStudy (anonymous):

add 1 1/4 + 1 1/5 and this will tell you how much pizza was eaten. After you get that answer, subtract it from 3 and that will tell you how much they took home.

OpenStudy (anonymous):

how to do the addition again, i forget

OpenStudy (anonymous):

don't understand

OpenStudy (anonymous):

1 1/4 = 5/20 1 1/5 = 4/20 ------------add 2 9/20 you can then turn 2 9/20 into an improper fraction... 2 9/20 = 49/20. now subtract 49/20 from 3

OpenStudy (anonymous):

is the answer 2 3/10

OpenStudy (anonymous):

3/1 - 49/20 = 60/20 - 49/20 = 11/20 check... 49/20 + 11/20 = 60/20 = 3 your answer is 11/20

OpenStudy (anonymous):

ok thanks , I think i understand now

OpenStudy (anonymous):

thats good....if you have any questions, just ask :)

OpenStudy (anonymous):

thanks